Two positive point charges 5.25 μC are fixed on the y-axis at y = 7.75 mm and y = -7.75 mm. A negative point charge -4.25 μC is located 10.0 mm on the +x-axis.

What is the x-component of the force on the -4.25 μC charge?

______N

What is the y-component of the force on the -4.25 μC charge?

_____N

What is the magnitude of the net force on the -4.25 μC charge if it is located at the origin?

______N

Where must the -4.25 μC charge be place to obtain the maximum force on it from the 2 charges q?

____mm

What is that maximum force (magnitude only)?

_____N
Answer #1

Fsino F YE7.7Smm 10mme ett Side P 2F COS9 y 7-75 mm FSino F Fi= F2 Fy FSine - FsinO Same 2fx -2 F COSS = 2 K9,92 Cos 2 7.75 tame 0 3777 17 7752+02 = 2x9 x103x 5-25X106 2.6S mm x1632 (12.6S XI3 1983. 93N Etx

F F= F-F d F S.u8mm Mm Hemo 7.75 548 Fnax 2F COSO 7 -12 2Xq Xlo 4.25X5.25 X10 7.7545-4X1032 2Sy. 2N Fmax